OpenStudy (anonymous):

the college paid $1858 for an order of 50 calculators. The store paid $9 for each scientific calculator and all graphing calculators cost the store $53. How many of each type was ordered?

OpenStudy (iurxos):

53*n+9*m=1858 and n+m= 50 where n is number of scientific and m is number of graphing caculators. Then n=50-m, which is fed to the first eqn to obtain: 53*(50-m)+9*m=1858 so -44m=1858-2650 and m=18. Then n=32